# Knowledge Graph Supported Benchmark and Video Captioning for Basketball

### 25 Jan 2024 v1

@@ 3.1 Multimodal Basketball Game Knowledge Graph
Table 1: Statistics of the labels in NBA. “succ.” and “fail.” are abbreviations of “success” and “failure”, respectively.
By analyzing the play-by-play data, game events can be divided into 9 categories, including “Foul”, “Rebound”, “Violation”, “Timeout”, “Freethrow”, “Enter game”, “Turnover”, “Jump ball”, and “Shot”. Each text description has its owns keyword, such as the description of the “Foul” event “Personal foul by G. Temple (drawn by A. Drummond)”, where the keyword is “foul”. Therefore, the type of this event is “Foul”. For this semi-structured data, the sentence is parsed into multi-tuples using the character index of the keyword and specific words that appear in the sentence, such as “drawn by”. To incorporate video information into the multi-tuples, the video timestamp is intended to be associated with the text description. OCR is used to identify the time on the scoreboard in each frame and record the timestamp of the current frame. Specifically, we employ the open-source OCR toolkit PaddleOCR [^3] https://github.com/PaddlePaddle/PaddleOCR and Tesseract-OCR [24] to recognize the time simultaneously. By using multiple OCR toolkits, we can improve the overall accuracy of recognition. This is because one toolkit may succeed where the other fails. In addition to event information, player and team information are also stored in multi-tuples. Through these efforts, a Multimodal Basketball Game Knowledge Graph containing 11489 events and 42870 relationships is constructed. An example of the Multimodal Basketball Game Knowledge Graph (MbgKG) is shown in Fig. 2. More details on the construction of MbgKG can be found in Appendix A.
3.2 Multimodal Basketball Video Game Captioning Dataset
To construct a dataset that contains videos, text descriptions, and external knowledge (i.e., player images and names), the pertinent data is extracted from MbgKG by utilizing the relations within MbgKG. As shot and rebound events are the most prevalent in basketball, we only extract the data about these events from MbgKG. To be more realistic, the text descriptions need to be further modified.
As shown in Fig. 4, the event-action relation in the graph is used to extract 9 kinds of events, and the video-time relation in the graph is used to obtain the video timestamp of the event. Specifically, we first roughly estimate the start and end timestamps of the event based on the existing timestamp. Each clip undergoes a manual review process to ensure accurate start and end timestamps. Subsequently, the video-description relation is used to obtain the text description, which is then matched with the corresponding video clip. We extract player-related information (player images and player names) from MbgKG through the team-player and player-name relations. Since each event involves certain players, only the involved individuals and their teammates need to be considered as candidate players for video annotation.
Using the sentence “B. Ingram misses 2-pt jump shot from 19 ft and defensive rebound by J. Winslow” as an example, we will explain our text modification process. Due to the difficulty in generating distance in video captioning task, the “from 19 ft” in the sentence is removed. Based on the attribute information of the player in MbgKG, the abbreviated name (“J. Winslow”) in the sentence is replaced with the full name (“Justise Winslow”). This change might increase the complexity of name generation but is more aligned with real-world scenarios. To improve text fluency, alterations are made based on the character index. Specifically, we identify the index of keywords like ”defensive” and apply a rule (“defensive/offensive rebound by SOMEONE” $\Longrightarrow$ “SOMEONE gets the defensive/offensive rebound”) to revise the sentence. Finally, there are a total 3977 videos [^4] The scale of the current dataset is not large enough, which is emergent to be enriched in our subsequent research., each of which belongs to one of 9 kinds of events. Each video has one text description and several candidate players information (images and names). We randomly select 3162 clips for training and 786 clips for testing. Tab. 1 shows the sample distributions across different labels of events. More details on the construction of MbgVC can be found in Appendix B.
In Tab. 2, the comparison of our dataset with MSR-VTT, YouCook and ActivityNet Captions further demonstrates the fine-grained details of our captioning annotations. MbgVC has the most sentences per second of 0.182, while the other datasets are all below 0.1, this reflects that our dataset has more detailed information in descriptions. Moreover, MbgVC has 1.73 verbs in a sentence on average, higher than 1.41 for ActivityNet Captioning and 1.37 for MSR-VTT. Similarly, the verb ratio of MbgVC which is computed by dividing the total number of verbs by the total number of words in the sentence is also much higher than other 3 datasets. This highlights that our dataset is primarily focused on the fine-grained actions of the subjects, aligning with our original intent.
3.3 Evaluation Metric
Existing video captioning benchmarks mainly adopt n-gram based metrics, including CIDEr [28], METEOR [1], Rouge-L [11], and BLEU [18]. BLEU evaluates the quality of the generated text by calculating the n-gram overlap between the generated text and the reference text. ROUGE-L pays more attention to generating the longest common subsequence between the generated text and the reference text, and evaluates sentence quality in this way. METEOR takes more factors into account, including stem matching, synonym matching, and word order. CIDEr uses TF-IDF [22] to assign different weights to n-grams of different lengths, then calculates the cosine similarities of n-grams between generated texts and reference texts and averages them to get the final score. These metrics primarily emphasize the consistency of text rather than the accuracy of semantics. In other words, if the generated text differs from the reference in structure or vocabulary, it may be penalized by these n-gram based evaluation metrics, even if the conveyed information is completely correct. However, for entity names, they cannot be replaced with the similar words.
To take into account the predictive accuracy of specific entity names, the entity name prediction accuracy $e_{acc}$ is introduced to accurately evaluate the accuracy of name prediction. $e_{acc}$ is a score calculated by dividing the number of correctly predicted names by the total number of names. Finally, we settle on a new metric, Game Description Score (GDS), as follows:
$gds=\frac{4\cdot e_{acc}+2\cdot cs+2\cdot ms+1\cdot rs}{9}\times 100,$ (1)
where $cs$, $ms$, and $rs$ refer to entity name prediction accuracy score, CIDEr score, METEOR score, and Rouge-L score, respectively. Since CIDEr involves the partitioning and matching of n-grams, BLEU is not included in the calculation. More detailed descriptions of GDS can be found in Appendix C.
@@ 3.3 Evaluation Metric
MbgVC 0.182 1.73 14.8%
Table 2: Comparisons of different video caption datasets. Numbers in bold indicate the highest value.
4 Knowledge Graph Supported Basketball Video Captioning
4.1 Problem Formulation
To generate precise and concise textual descriptions of the basketball live text broadcast, we introduce the concept of a Knowledge-Aided Basketball Video Captioning (KABVC) task. In this task, the model is required to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the external knowledge. This task can be formulated as: given the basketball video $V_{b}$, the objective is to select video-related player knowledge $K_{p}=s\left(V_{b}\right)$ and generate the text description $D_{v,k}$.
${\rm KABVC}:D_{v,k}=m\left(V_{b},K_{p}\right)=m\left(V_{b},s\left(V_{b}\right)\right),$ (2)
@@ 4.1 Problem Formulation
Model CIDEr METEOR Rouge-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 GDS
V2C 13.7 18.7 45.9 50.1 39.5 27.0 14.9 12.3
Clip4Caption 19.4 26.7 51.2 49.1 42.5 35.4 28.8 25.0
SwinBert 20.1 26.5 49.0 47.8 41.4 34.5 28.4 25.6
EAC 138.5 28.0 54.9 53.1 46.4 38.8 32.4 56.9
Table 3: Combined inference performance on MbgVC. Numbers in bold indicate the best performance.
4.2 Proposed Model
For the ${\rm KABVC}$ task, we propose a model with an encoder-decoder framework named Entity-Aware Captioner (EAC). The overall structure of EAC framework is shown in Fig. 4. Given the raw video frames which are of size $T\times 3\times H\times W$, consisting of $T$ frames and each has $3\times H\times W$ pixels. We feed them into Resnet-18 [7], which is the video encoder in EAC, and extract the video feature $F_{T}\in{\rm R}^{T\times D_{r}\times h\times w}$. $D_{r}$ is the hidden size of Resnet-18. The global representations of $F_{T}$ are then fed into Bi-GRU module to further encoder temporal information and obtain the feature $F_{v}$.
@@ 4.2 Proposed Model
where $W_{1}\in{\rm R}^{D_{r}\times D}$ is the linear mapping layer, which maps the video feature to text space. $D$ is the hidden size of the decoder module. ${\rm GRU}\left(\cdot\right)$ denotes the Bi-GRU module and ${\rm AvgPool}\left(\cdot\right)$ denotes the average pooling layer.
Each video has the corresponding candidate $N$ players’ information as the external knowledge to assist the generation of text with the specific entity names. Each player image is of size $3\times H_{p}\times W_{p}$. For candidate players’ images, Resnet-18 is used to extract the features $F_{N}\in{\rm R}^{N\times D_{r}\times h_{p}\times w_{p}}$. The global features $F_{p}\in{\rm R}^{N\times D_{r}}$ of $N$ images are obtained by Eq. 4.
@@
where $W_{2}\in{\rm R}^{D_{r}\times D}$ is the linear mapping layer, which maps the image feature to text space.
The text encoder of large language model T5 [21] is used to transform candidate names into a sequence of embeddings $T_{n}\in{\rm R}^{N\times D}$. After that, the image features $F_{p}$ and name features $T_{n}$ are added by corresponding positions to obtain multimodal player features $F_{m}\in{\rm R}^{N\times D}$. Basketball events involve interactions between multiple players. For example, in “Myles Turner makes the 2pt layup with an assist from Tyrese Haliburton”, “Myles Turner” scores a two-point layup with the assist of “Tyrese Haliburton”. Therefore, the multimodal player features $F_{m}$ are fed into the multi-head self-attention module to model the relationship among candidate players. The player features based on attention and the video features containing motion information are then concatenated and fed into the decoder ${\rm\Psi_{T5}}\left(\cdot\right)$ of T5 to obtain the text description $C$.
$C={\rm\Psi_{T5}}\left(W_{3}\left({\rm Concat}\left[{\rm H_{M-ATT}}\left(F_{m}\right),F_{v}\right]\right)\right),$ (5)
where $W_{3}\in{\rm R}^{D\times D}$ is the linear mapping layer, which maps the concatenated feature to the vector space of the T5 model. ${\rm Concat}\left[,\right]$ denotes the Concatenation function in Python, and ${\rm H_{M-ATT}\left(\cdot\right)}$ denotes the multi-head self-attention module.
The Bi-GRU module, 3 linear layers, multi-head self-attention and text decoder are trained to maximize the log-likelihood over the training set given by Eq. 6.
$L_{\Theta}=\sum_{t=1}^{N_{C}}log{\rm P_{r}}\left(y_{t}|y_{t-1},\left[F_{m},F_{v}\right];\Theta\right),$ (6)
@@ 5 Experiments
To verify the performance of the proposed model, EAC is compared with the advanced video captioning models. We further conduct the ablation experiments to verify the effectiveness of each conponent in EAC.
5.1 Implementation Details
@@
The proposed EAC uses ResNet-18 pretrained on the ImageNet dataset [23] as the visual encoder, which hidden size $D_{r}$ is 512. EAC is trained on the proposed MbgVC dataset with 100 training epochs. In MbgVC dataset, $T$ frames are sampled by using segment-based method [29]. Each frame size is $720\times 1280$. Concretely, $T$ is set to 18. The size of player’s image is $120\times 180$ and the number of candidate players in each video is not fixed. For the large language model T5, the hidden size is 768. It is worth noting that the parameters of the visual encoder and text encoder are frozen during training. During the training stage, the model is optimized by ADAM [8] with the learning rate of 3e-5 and weight decay of 1e-4. Beam search with a beam size of 2 is used for inference. The proposed EAC is implemented with Python 3.9 and PyTorch 1.12, and is performed on a server with an Nvidia 3090ti GPU.
5.2 Main Results
@@
EAC is compared with three advanced video captioning models, including V2C [5], Clip4Caption [26], and SwinBert [12]. V2C is the Transformer-based model that generates relevant commonsense descriptions of the given video. Clip4Caption employs CLIP to acquire the aligned visual-text representation for better generating the text descriptions. SwinBert introduces Video Swin Transformer to encode spatial-temporal representations from video frames.
@@
As shown in Tab. 3, EAC outperforms other 3 models by a large margin on CIDEr. This notable performance can be attributed to the way CIDEr calculates cosine similarities of n-grams between the generated text and reference text. Highly accurate name prediction contributes to a higher cosine similarity between the n-grams of the generated text and the reference text. V2C, Clip4Caption, and SwinBert exhibit lower performance in accurately generating names, resulting in much lower CIDEr scores. Although EAC achieves the best performance across all metrics, the differences are not as pronounced in METEOR, Rouge-L, and BLEU. This can be attributed to the concise nature of the text descriptions in the dataset, primarily consisting of names and fine-grained actions. Notably, EAC outperforms SwinBert 31.3% on GDS and outperforms Clip4Caption 31.9% on GDS. The higher GDS indicates better overall model performance, particularly in name prediction. These results underscore the capability of our model to generate accurate names and fine-grained actions in live text broadcast task. Qualitative results can be found in Appendix D.
Model Ki Kn Attn. Bi-GRU CIDEr METEOR Rouge-L BLEU-4 GDS
20.1 20.6 40.9 23.0 19.1
18.0 20.0 38.0 23.2 16.3
110.6 25.2 49.1 27.3 46.7
115.7 25.6 53.1 28.1 49.8
119.3 26.8 53.3 29.5 51.4
122.2 27.3 54.6 30.5 52.5
138.5 28.0 54.9 32.4 56.9

### 28 Feb 2024 v2

@@ III-B Multimodal Basketball Video Game Captioning Dataset
TABLE III: Comparisons of different video caption datasets. Numbers in bold indicate the highest value.
@@
VC_NBA_2022 0.182 1.73 14.8%
IV Knowledge-based Entity-aware Basketball Video Captioning
IV-A Problem Formulation
To generate precise and concise text descriptions of the basketball live text broadcast, we introduce the concept of KEBVC task. In this task, the model is required to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the additional knowledge. This task can be formulated as: given the basketball video $V_{b}$, the objective is to select video-related player knowledge $K_{p}=s\left(V_{b}\right)$ and generate the text description $D_{v,k}$.
${\rm KEBVC}:D_{v,k}=m\left(V_{b},K_{p}\right)=m\left(V_{b},s\left(V_{b}\right)\right),$ (1)
@@ IV-A Problem Formulation
TABLE IV: Combined inference performance on VC_NBA_2022. Numbers in bold indicate the best performance.
Model CIDEr METEOR Rouge-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 RoleF1
V2C 13.7 18.7 45.9 50.1 39.5 27.0 14.9 0.0
Clip4Caption 70.4 26.7 51.2 49.1 42.5 35.4 28.8 8.3
SwinBert 69.1 26.5 49.0 47.8 41.4 34.5 28.4 7.4
KEANet 138.5 28.0 54.9 53.1 46.4 38.8 32.4 20.6
Figure 7: Detailed architecture of the entity-aware module.
IV-B Proposed Model
For the ${\rm KEBVC}$ task, we propose a knowledge guided entity-aware video captioning network (KEANet) based on a candidate player list in encoder-decoder form. The overall structure of KEANet framework is shown in Fig. 6. Given the raw video frames which are of size $T\times 3\times H\times W$, consisting of $T$ frames and each has $3\times H\times W$ pixels, we feed them into the CNN-based visual encoder in KEANet and extract the video feature $F_{T}\in{\rm\mathbb{R}}^{T\times D_{r}\times h\times w}$. $D_{r}$ is the hidden size of visual feature. The global representations of $F_{T}$ are then fed into Bi-GRU module to further encode temporal contextual information and obtain the feature $F_{v}\in{\rm\mathbb{R}}^{T\times D}$.
@@ IV-B Proposed Model
where $W_{1}\in{\rm\mathbb{R}}^{D_{r}\times D}$ is the linear mapping layer, which maps the video feature to text space. $D$ is the hidden size of the decoder module. ${\rm GRU}\left(\cdot\right)$ denotes the Bi-GRU module and ${\rm AvgPool}\left(\cdot\right)$ denotes the average pooling layer.
Each video has the corresponding candidate $N$ players’ information as the additional knowledge to assist the generation of text with the specific entity names. Each player image is of size $3\times H_{p}\times W_{p}$. For candidate players’ images, we use the CNN-based visual encoder to extract the features $F_{N}\in{\rm\mathbb{R}}^{N\times D_{r}\times h_{p}\times w_{p}}$. The global features $F_{p}\in{\rm\mathbb{R}}^{N\times D}$ of $N$ images are obtained by (3).
@@
where $W_{2}\in{\mathbb{R}}^{D_{r}\times D}$ is the linear mapping layer, which maps the image feature to text space.
TABLE V: Combined inference performance on Goal. Numbers in bold indicate the best performance.
Method CIDEr METEOR Rouge-L BLEU-1
V2C 0.13 2.11 3.42 3.38
Clip4Caption 2.17 5.01 5.45 5.71
SwinBERT 2.20 5.12 5.32 5.65
KEANet 3.68 6.38 10.45 14.88
The text encoder of large language model T5 [43] is used to transform candidate names into a sequence of embeddings $T_{n}\in{\mathbb{R}}^{N\times D}$. After that, the image features $F_{p}$ and name features $T_{n}$ are added by corresponding positions to obtain multimodal player features $F_{m}\in{\mathbb{R}}^{N\times D}$. Basketball events involve interactions between multiple players. For example, in “Myles Turner makes the 2pt layup with an assist from Tyrese Haliburton”, “Myles Turner” scores a two-point layup with the assist of “Tyrese Haliburton”. Therefore, we design the entity-aware module to highlight key players from a candidate player list. The architecture of entity-aware module is shown in Fig. 7. entity-aware module first uses the cross-attention to fuse player features and video features to obtain $F_{f}$. This connects the players to the video content, putting the players in a specific scene. $F_{a}$ is sent to the multi-head self-attention sub-module to model the relationship among candidate players. The player features based on attention and the video feature are then concatenated and fed into the decoder ${\rm\Psi_{T5}}\left(\cdot\right)$ of T5 to obtain the text description $C$.
$C={\rm\Psi_{T5}}\left(W_{3}\left({\rm Concat}\left[{\rm E_{a}}\left(F_{m},F_{v}\right),F_{v}\right]\right)\right),$ (4)
where $W_{3}\in{\mathbb{R}}^{D\times D}$ is the linear mapping layer, which maps the concatenated feature to the vector space of the T5 model. ${\rm Concat}\left[,\right]$ denotes the concatenation function in Python, and ${\rm E_{a}\left(\cdot\right)}$ denotes the entity-aware module.
The Bi-GRU module, 3 linear layers, entity-aware module and text decoder are trained to maximize the log-likelihood over the training set given by (5).
$\mathcal{L}_{\Theta}=\sum_{t=1}^{N_{C}}log{\rm P_{r}}\left(y_{t}|y_{t-1},\left[F_{m},F_{v}\right];\Theta\right),$ (5)
@@ V Experiments
To verify the performance of the proposed model, KEANet is compared with the advanced video captioning models on VC_NBA_2022 and Goal [19]. We further conduct the ablation experiments to verify the effectiveness of each conponent in KEANet.
V-A Implementation Details
@@
The proposed KEANet uses ResNet-18 [44] pretrained on the ImageNet dataset [45] as the visual encoder, which hidden size $D_{r}$ is 512. KEANet is trained on the proposed VC_NBA_2022 and Goal datasets with 100 training epochs. In these datasets, $T$ frames are sampled by using segment-based method [46]. Each frame size is $1280\times 720$. A high resolution of video frame is helpful for the model to understand the players’ action in the video. Concretely, $T$ is set to 18. The size of player’s image is $180\times 120$ and the number of candidate players in each video is not fixed. For the large language model T5, the hidden size is 768. It is worth noting that the parameters of the visual encoder and text encoder are frozen during training. During the training stage, the model is optimized by ADAM [47] with the learning rate of 3e-5 and weight decay of 1e-4. Beam search with a beam size of 2 is used for inference. The proposed KEANet is implemented with Python 3.9 and PyTorch 1.12, and is performed on a server with an Nvidia 3090ti GPU.
Figure 8: Qualitative results on VC_NBA_2022 dataset. (V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; KEA: our proposed model; GT: the ground truth). The different specific entity names are marked in red and blue, respectively. And the fine-grained actions are marked green. Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names. So names are replaced with the special token $<$UNK$>$.
V-B Evaluation metrics
Existing video captioning benchmarks mainly adopt ngram-based metrics, including CIDEr [48], METEOR [49], Rouge-L [50], and BLEU [51]. BLEU evaluates the quality of the generated text by calculating the n-gram overlap between the generated text and the reference text. Rouge-L pays more attention to generating the longest common subsequence between the generated text and the reference text, and evaluates sentence quality in this way. METEOR takes more factors into account, including stem matching, synonym matching, and word order. CIDEr uses TF-IDF [52] to assign different weights to n-grams of different lengths, then calculates the cosine similarities of n-grams between generated texts and reference texts and averages them to get the final score. These metrics primarily emphasize the consistency of text rather than the accuracy of semantics. In other words, if the generated text differs from the reference in structure or vocabulary, it may be penalized by these n-gram based evaluation metrics, even if the conveyed information is completely correct. However, for entity names, they cannot be replaced with the similar words. To accurately evaluate the performance of specific entity name generation, we introduce RoleF1 score (F1 score of entity name generation) from the work [39]. For RoleF1, we extract entity names from the ground truth and the generated texts. Recall is used to measure how the generated texts covers the entities appearing in the game video clip. Taking into account that these generated entity names could also come from the model’s hallucinations, such as from incorrect games, we will also consider Precision. Finally, RoleF1 score is calculated with Precision and Recall.
Figure 9: The relationship between metric CIDEr and RoleF1. C4C denotes Clip4Caption.
V-C Results on VC_NBA_2022
@@
KEANet is compared with three advanced video captioning models, including V2C [37], Clip4Caption [32], and SwinBert [7]. V2C is the Transformer-based model that generates relevant commonsense descriptions of the given video. Clip4Caption employs CLIP to acquire the aligned visual-text representation for better generating the text descriptions. SwinBert introduces Video Swin Transformer to encode spatial-temporal representations from video frames. V2C, Clip4Caption, and SwinBert take only videos as input, and are all trained and tested on our proposed dataset.
@@
As shown in Table IV, KEANet outperforms other 3 models by a large margin on CIDEr. This notable performance can be attributed to the way CIDEr calculates cosine similarities of n-grams between the generated text and reference text. Highly accurate name prediction contributes to a higher cosine similarity between the n-grams of the generated text and the reference text. V2C, Clip4Caption, and SwinBert exhibit lower performance in accurately generating names, resulting in much lower CIDEr scores. Although KEANet achieves the best performance across all metrics, the differences are not as pronounced in METEOR, Rouge-L, and BLEU. This can be attributed to the concise nature of the text descriptions in the dataset, primarily consisting of names and fine-grained actions. Notably, KEANet outperforms SwinBert 13.2% on RoleF1 and outperforms Clip4Caption 12.3% on RoleF1. The higher RoleF1 indicates better name prediction. These results underscore the capability of our model to generate accurate names and fine-grained actions in live text broadcast task.
Fig. 8 shows the qualitative results on VC_NBA_2022 dataset, including generation results of V2C, Clip4Caption, SwinBert and our KEANet model. V2C, Clip4Caption and SwinBert can correctly generate some actions but fail to generate correct entity names because these names never appear during training. However, with the help of the additional knowledge, our KEANet could well relate entities in video clips with the fine-grained actions. These cases demonstrate that our newly proposed KEANet model more consistently with the requirement of practical application.
Further, we study the relation between the metric CIDEr and RoleF1. To this end, we plot the performance of KEANet and Clip4Caption with different RoleF1 scores in Fig. 9. The plot shows that this relationship is approximately linear: more accurate entity name generation imply better model performance.
V-D Results on Goal
Goal is a benchmark which contains over 8.9k soccer video clips, 22k sentences and 42k knowledge triples. These sentences were converted from the commentator’s audios. This type of comment sentences have a certain degree of colloquialism and are relatively long. Therefore, this dataset poses certain challenges. On this basis, we filter out sentences and videos that do not contain entity names. We modify the format of the dataset to be the same as VC_NBA_2022. In addition, we save the names of teams and players in the dataset. The revised Goal dataset is shown in Fig. 10.
KEANet is compared with three advanced video captioning models, including V2C, Clip4Caption, and SwinBert. V2C, Clip4Caption, and SwinBert take only videos as input, and are all trained and tested on the revised Goal dataset. As shown in Table V, we compare the performance of V2C, Clip4Caption, SwinBERT, and KEANet on several metrics including CIDEr, METEOR, Rouge-L, and BLEU-1. KEANet is supported by additional knowledge that it can generate text descriptions with entity names. Therefore, the KEANet outperforms other models in all metrics.
Fig. 11 shows the qualitative results on Goal dataset, including generation results of V2C, Clip4Caption, SwinBert and our KEANet. Due to the lack of relevant tokenizer tools in V2C, it cannot process text with entity names and can only use special symbols to replace specific entity names. Compared to the other three models, KEANet can generate correct entity names and partial actions. The above results indicate that, even in challenging tasks such as sports commentary, with the support of additional knowledge, the model can generate text with entity names.